OpenStudy (anonymous):

Suppose the area of a square plaque is 169 square inches. What is the length of each side of the plaque?

jimthompson5910 (jim_thompson5910):

Area of a square = s^2 where s is the side length So A = s^2, but we know that A = 169, so 169 = s^2 Take the square root of both sides and flip the equation to get s = 13. So the length of each side of the square is 13 inches. Note: 13^2 = 169

OpenStudy (anonymous):

\[\sqrt169\]\[10\sqrt69\]

jimthompson5910 (jim_thompson5910):

not sure how you're getting that figure elizated

OpenStudy (anonymous):

well the area of a square is x(x) so the sqrt of the area would be a side length.

jimthompson5910 (jim_thompson5910):

exactly, but \[\large \sqrt{169}=\sqrt{13^2}=13\]

jimthompson5910 (jim_thompson5910):

since \[\large 13^2=169\]

OpenStudy (anonymous):

oh woops. sorry i didn't do this right. sorry. ignore my answer... jim is the right one. thanks for pointing that out.

OpenStudy (anonymous):

issok :)